4. For the experiment, Mg was reacted with HCI to produce H, Starting with a 0.0145g piece of Mg using (Mg 24gmol; H2 2.0g/mol) a) Write balanced chemical equation and determine theoretical yield of H
Answer #1

4 a)

Reaction:

Mg + 2HCl \rightarrow MgCl_{2} + H_{2}

As per the above equation 1 mole of Mg produces 1 mole of H2

# moles of Mg = mass of Mgl/Atomic wt of Mg = 0.0145/24 = 0.00060 moles = # moles of H2 produced

Mass of H2 = moles * mol wt of H2 = 0.00060 * 2 = 0.0012g

Theoretical yield = 0.0012 g

b) P_{total} = P_{O_{2}} + P_{H_{2}O}

P_{O_{2}} = P_{total} - P_{H_{2}O}

P(O2) = 764 - 22 = 742 atm ( 1 mmHg = 1 torr)

Use PV= nRT

PV = m(H2) R T/mol wt of H2

mass of H2 = P * V * mol wt of H2/R * T = 742 * 0.021 * 2/62.40 * 298 = 0.0017g